my problem is realy not a question,its a statement. it says,given the circle is inscribed in the square,find the area of the circle. on the outside right of the square it has 28ft.

Answer by TimothyLamb(4379)

---
this probably means that each side of the square is 28 ft, if so here's your answer:
---
r = radius of circle = 28/2 = 14 ft
a = area of circle = pi*r^2
---
a = pi*14^2
a ~= 615.752 sq.ft
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php